# Scaling the Authoring of AutoTutors with Large Language Models

### v2

@@ 1. Introduction
The design process of AutoTutor and other similar intelligent tutoring systems can be split into two distinct stages, carried out by different, yet potentially overlapping, sets of stakeholders. The first stage is the general framework level design, carried out mostly by learning scientists (possibly along with computer scientists). This includes designing the pedagogy, the mechanisms required to interact with the student, and potentially even authoring tools. This is typically done only once for an entire domain. The second stage, performed by either teachers or domain experts, is to author the scripts111see Appendix A in (Graesser et al., 2000) for an example script, which needs to be done separately for every topic, question, or concept. In the standard framework, both these stages are done by humans. In contrast, in the LLM-only framework discussed in the previous paragraph, both these stages are carried out (implicitly or explicitly) by LLMs.
In this paper, we consider a mixed approach where we only delegate the job of domain experts to an LLM, while the first stage is still left to learning experts. This idea has previously been hinted at by Cai et al. (Cai et al., 2019), who propose using a question-answering system to automatically list potential misconceptions for a given topic. We extend this idea to have an LLM author the entire script for the tutoring system, which is accurate enough to be deployed without human validation but is fully open to human validation.
Our contributions in this paper are as follows:
We adapt the Dialog Based tutoring framework used by ITS’s to make it possible for a Large Language Model to author it.
We design MWPTutor, a tutor for math word problems based on the above principles.
We compare the performance of MWPTutor with that of GPT4 being used as a tutor both quantitatively and qualitatively.
@@ 2.1. Rule-Based Systems
To combat the above scarcity, Intelligent Tutoring Systems (ITSs) were already being developed since the early 1980s (Sleeman and Brown, 1982). The idea was to roleplay a human tutor with a computer program that could interact with the student and fulfill the requirement of personalized tutoring. The computer program may not fully replace human tutors but often provides an affordable and scalable addition. An important aspect that differentiates ITSs from simply showing the students the problem solution is their ability to provide step-by-step feedback which promotes long-term learning as opposed to short-term problem-solving success. This was often implemented using rules that check for certain errors using string-matching and act on them with pre-defined templates. For example, LISP Tutor (Anderson et al., 1985) contains rules to account for common programming errors made by students.
@@
With concepts and tutoring strategies getting more complex, writing the rules governing the finite state transducers quickly becomes time-consuming. Not knowing the student misconceptions in advance, domain experts have to come up with rules for numerous possibilities, introducing a different set of scaling issues. One hour of interaction with a student costs over 100 hours of domain expert time to author the ITS (Koedinger et al., 2004). This is in addition to the time spent by learning scientists designing the pedagogy and dialog strategies. However, this is usually done only once for the topics in a subject area. As a result, this limits the AutoTutor to a small set of subject areas, restricting their impact. Nevertheless, there are many works in the space of rule-based ITS (Paladines and Ramirez, 2020), which have been reasonably successful in their domains.
@@ 3. Design Principles
The optimal tutoring support differs between the two examples. While the math word problem starts with subquestions to compute intermediate quantities, answering the history questions first requires providing historical context.
@@ 3.3. Scaffolding the LLM
While modern LLMs fine-tuned with RLHF are quite good at following instructions, they can at times struggle with more complex tasks. Because of the probabilistic nature of LLMs, reattempting the same problem can lead to different answers. To take the best advantage of these facts, we want to have automated checks for every LLM turn in the strategy space, so calls can be repeated until some checks have been met. For example, D3 could require that the numerical answer (450 for A1) is included in the utterance, while P2 could require the missing keywords to be absent. These checks can take different forms depending on the domain, like limiting the length of an utterance or avoiding the use of more complex language.
3.4. Control and Reliability
One of the primary goals of having LLMs author a state space rather than having them interact directly with the students is that in this way, one can examine and trust the model without having to necessarily trust the LLM used to generate it. Based on this, we store and make it possible to inspect beforehand the entire utterance space. Changes to the utterances themselves can be made on-the-go, e.g. if a teacher feels it is better to split our example in P2 into two sentences, they can simply edit it to say “Who was assassinated?” and “Who carried out the assassination” with no further changes required. Changes to the solution space, however, need the utterances to be regenerated.
Caching the utterance LLMs like GPT also makes the system immune to changes made to GPT in terms of availability, cost, or rate limits. For open LLMs, this allows for the use of Larger LLMs which not every user can run locally. The most important benefit, however, is the potential for teachers/domain experts to edit the state space should they want, therefore allowing us to combine human expertise with machine efficiency. Live calls to an LLM can often help and can be incorporated into the state space. For example, in the MWP we could make a call to GPT to take over a step in between D2 and D3 if we want to minimize the chance of D3. Should live calls be implemented, however, there should ideally be non-LLM fallbacks ie we should have cached and/or procedural alternatives available should a call to an LLM not be possible at deployment time.
@@ 4. MWPTutor
We now detail the design process of our system MWPTutor, an instantiation of the above principles. MWPTutor is adapted to help students solve math word problems. The scenario, originally defined by MathDial (Macina et al., 2023), involves a student approaching a tutor after unsuccessfully attempting to solve a math word problem. The student’s incorrect solution is provided as a part of the input. GSM8k (Cobbe et al., 2021), the source for MathDial’s problems includes a single step-by-step solution to each problem, which can be seen as a fulfillment of the solution step space. However, to not become constrained to the availability of such solutions, and to allow diverse solution paths, we restrict MWPTutor to only the final numerical answer to the problem, and generate the solutions ourselves.
@@ 4.3. Pedagogical Strategy
| Pump |
asks the student to propose the next step. If the student proposes the step correctly, we ask them to solve it. |
| Hint |
is only triggered if the student fails to propose a valid next step, or fails to execute the step they propose. It essentially gives out the step and asks the student for the calculation. This utterance should take the form of a question, and the only way to correctly answer the question should be to successfully execute the step. |
| Prompt |
is triggered if the student fails to execute the step after the hint. This should push the student closer to the answer, without actually revealing the answer. |
| Assertion |
is triggered if the student fails to answer even after the prompt. This should simply give out the answer, and have the student acknowledge it. |
In our running example, hints, prompts, and assertions would correspond to D1, D2, and D3 respectively. The pump is not included.
@@ 4.4. Answer and Path Detection
To execute the pedagogical strategies described above, we need to detect if the student has given a correct answer. Further, the hint state should only trigger if the student fails to propose a correct follow-up. For this, we need to have a way to match paths. We run into the same matching problems described in Section 4.2, and resort to similar basic techniques. The correct answer detection algorithm is shown in Algorithm 2. For states other than pump, it takes a single reference step and checks if its RHS is present in the student utterance. For the pump state, all possible followups of the last step are passed as reference steps. If it can find the RHS of any of these in the student utterance, it marks that as the next step and proceeds to the next step. If not, it considers a path match to have occurred if every operand on the left-hand side of some step is present. If multiple steps fulfill this, or no step does, we pick the one with the best BertScore alignment to the student utterance.
The fact that we currently ignore all non-numeric content in the student response is one of the major limitations in our work, as it can miss a large number of “correct” utterances from the student, but we defer the resolution of this to future work.
@@ 4.5. Putting it All Together
Figure 3 summarises the overall state space of MWPTutor. It can be seen as a simplified version of the state space of AutoTutor (see Figure 1 in (Cai et al., 2019)). The expectation matching is implicitly carried out by the solution aligner and the pump state. To obtain to full finite state transducer structure, one must roll out the states along the solution space. Excluding the unlikely case where the student fails to repeat the answer after the assertion state, the overall conversation should not last more than utterance pairs, where is the number of steps.
@@ 5.1. Setup
At the time of testing, GPT4555gpt-4-1106-preview queried in January 2024 was considered to be one of the best-performing LLMs, so we use it both as a backend for MWPTutor and as the baseline single prompt LLM. To best align the single prompt GPT4 with the MWPTutor doctrine, we ask it to roleplay a tutor, setting the primary objective of getting to the correct answer as the goal of the conversation, and the secondary objective of not revealing the answer as a personality trait. Further, to test the potential of an open-source LLM as a backend, we have a second version of MWPTutor with a MetaMath-13B (Yu et al., 2023). MetaMath is a version of Meta AI’s open source LLM LLaMa666https://llama.meta.com/ fine-tuned to solve math word problems, and is, to the best of our knowledge, the best open LLM for solving math word problems. We refer to these three models as MWPTutor, GPT4, and MWPTutor respectively.
@@ 5.2. Conversation Evaluation
We evaluate both success and telling for the MultiArith dataset by directly looking for the answer in student and tutor utterances respectively, and calculate the overall scores based on that. The top half of Table 1 summarises our results. All models are able to transfer their near-perfect solving scores to near-perfect tutoring scores. GPT4 performs slightly worse, although it has a telling rate. Both versions of MWPTutor have a perfect success rate. It must be noted that the metric is noisy since it only requires the number to match, with no restriction on the context, e.g. in the problem “How many groups of can be made out of boys and girls?” if the tutor mentions there are boys, we would count it as telling as the final answer happens to be . This means that the real rates could be somewhat lower for all models, and the difference between them is not significant. Nevertheless, such errors are unlikely to be frequent, and all models work reasonably well on these problems.
| MWPTutor | GPT4 | |||
| Metric | LLaMa | GPT | ||
MultiArith |
Success Rate | 1.00 | 1.00 | 0.98 |
| Telling Rate | 0.01 | 0.01 | 0.00 | |
| Adjusted Success | 0.99 | 0.99 | 0.98 | |
| Tutoring Score | 0.99 | 0.99 | 0.98 | |
MathDial |
Success Rate | 0.91 | 0.96 | 0.82 |
| Telling Rate | 0.33 | 0.11 | 0.00 | |
| Adjusted Success | 0.58 | 0.84 | 0.82 | |
| Tutoring Score | 0.71 | 0.90 | 0.82 | |
@@
Table 1 summarises our result. While GPT4 maintains its 0% telling rate, MWPTutor gets more problems correct overall, and more when adjusted for telling. This gives MWPTutor an advantage on GPT4 in terms of the final score. MWPTutor also gets a higher success rate than GPT4, which is somewhat surprising since its backend model is significantly weaker than GPT4 in the MWP Solver setting (Yu et al., 2023). This increased success rate however comes at a cost of a very high telling rate, giving it a significantly poorer tutoring score overall, so we do not recommend the use of MetaMath as the backend model unless GPT4-like models need to be avoided for some external reason. One thing worth mentioning here is that once problematic questions have been identified for either version of MWPTutor, one can have human annotators redo only those questions, thereby increasing the success rate, something that is not possible for GPT4.
@@ 5.3. Utterance Evaluation
A major shortcoming of AI-on-AI-tutoring is that the AI student may not be affected by the sometimes subpar tutor responses. This is more common in MWPTutor than in GPT4, as the former can only pick from a pre-defined set of responses and not create new utterances on the fly in response to finer details in the student requirements. To understand if this is a major issue, we hired human evaluators who worked in the education sector and asked them to evaluate utterances made by MWPTutor for MathDial. Because the pumps are generic and very similar to each other and assertions essentially boil down to reading out the answers, we focus only on the hints and prompts in our evaluation.
@@
Across the conversations, MWPTutor and the student model went over 135 solution steps. Of these went to the hint state, while reached the prompt state. We presented each of these utterances to human evaluators, with the past conversation, the original math word problem, and the current solution step as context. We boil down the requirements set on hints and prompts set in Section 4.3 to the following binary criteria which we asked the evaluators to evaluate the utterances on:
Grammatical: The utterance is linguistically and grammatically correct.
Not Reveal: The utterance does not reveal the final answer for the step.
@@
Relevant (hints only): The hint clearly addresses the step, ie, the only way to correctly answer the question is to complete the current step.
@@
Not Redundant (hints only): The hint is not redundant to what the student already said in the previous step ie it makes sense to ask this question after the previous response.
@@
Specific (prompts only): The prompt is specific to the step, and not just a vague lead-on statement.
@@
Helpful (prompts only): The prompt can potentially help the student make progress in the current step, given the previous student’s utterance.
@@
The annotators were paid the Prolific suggested wage of 10 GBP/hour. We show the annotation user interface in Figure 6 and some example annotations in Table 2 and Table 3. Figures 4 and 5 summarise the responses from the evaluators. Almost all utterances are considered Grammatical by the evaluators, which is not surprising given they were generated by GPT4. Most of the utterances are also marked to be non-revealing, which is also not surprising given that any proposed utterances that directly revealed the answer were assigned to be regenerated. The evaluators further seem to agree that most hints were relevant to the step and that most prompts are specific, implying that they do what is appropriate given the current solution step.
@@
The biggest potential weakness of MWPTutor comes in the form of its utterances making sense in the context of previous student utterances due to issues discussed in Section 4.4, which is where the final two metrics come in. In this regard, all prompts are considered to be helpful by a majority of evaluators which is a great sign. Hints, however, are found to be redundant in almost a third of cases, showing that handling of student’s responses to pumps could be better. The current heuristic of matching the left-hand side is too demanding, and highly likely to ignore the student even if they propose a valid next step to the problem, which can break student immersion. Note that of the steps did not proceed to the hint state, implying more than half of the correct proposals by the student are being identified as such. Further, in a vast majority of cases where a correct proposal is misidentified, there is disagreement between human evaluators indicating there might be some merit in the rejection, though this is also the case for a majority of correct rejections.
@@
Overall these evaluations show that utterances generated for MWPTutor by the GPT4 backend are consistent with the strategy definitions, and the transition function uses them reasonably well.

### v3

@@ 1. Introduction
The design process of AutoTutor and other similar intelligent tutoring systems can be split into two distinct stages, carried out by different, yet potentially overlapping, sets of stakeholders. The first stage is the general framework level design, carried out mostly by learning scientists (possibly along with computer scientists). This includes designing the pedagogy, the mechanisms required to interact with the student, and potentially even authoring tools. This is typically done only once for an entire domain. The second stage, performed by either teachers or domain experts, is to author the scripts111See Appendix A in (Graesser et al., 2000) for an example script., which needs to be done separately for every topic, question, or concept. In the standard framework, both these stages are done by humans. In contrast, in the LLM-only framework discussed in the previous paragraph, both these stages are carried out (implicitly or explicitly) by LLMs.
In this paper, we develop a mixed approach where we only delegate the job of domain experts to an LLM, while the first stage is still left to learning experts. This idea has previously been hinted at by Cai et al. (Cai et al., 2019), who propose using a question-answering system to automatically list potential misconceptions for a given topic. We extend this idea to have an LLM author the entire script for the tutoring system. This authoring process includes internal checks and guardrails, making the results accurate enough to be deployed without human validation. However, should validation be essential, we also describe a version that can be checked by humans to increase trust.
In our experiments, we find that our system is better at tutoring an AI-simulated student than instructed, but otherwise free-form GPT4. We also find that human evaluators judge that individual utterances are indeed following the specified pedagogy. However, manual analysis of the conversations shows that there is still some scope for improvement. In summary, our contributions in this paper are as follows:
We adapt the Dialog Based tutoring framework used by ITS’s to make it possible for an LLM to author it.
We design MWPTutor, a tutor for math word problems which can function both with or without live calls to an LLM based on the above principles.
We compare the performance of both versions of MWPTutor with that of GPT4 being used as a tutor both quantitatively and qualitatively.
@@ 2.1. Rule-Based Systems
To combat the above scarcity, Intelligent Tutoring Systems (ITSs) were already being developed since the early 1980s (Sleeman and Brown, 1982). The idea was to roleplay a human tutor with a computer program that could interact with the student and fulfill the requirement of personalized tutoring. This may not fully replace human tutors but often provides an affordable and scalable addition. An important aspect that differentiates ITSs from simply showing the students the problem solution is their ability to provide step-by-step feedback which promotes long-term learning as opposed to short-term problem-solving success. This was often implemented using rules that check for certain errors using string-matching and act on them with pre-defined templates. For example, LISP Tutor (Anderson et al., 1985) uses rules to account for common programming errors made by students.
@@
With concepts and tutoring strategies getting more complex, writing the rules governing the finite state transducers quickly becomes time-consuming. Not knowing the student misconceptions in advance, domain experts have to come up with rules for numerous possibilities, introducing a different set of scaling issues. One hour of interaction with a student costs over hours of domain expert time to author the ITS (Koedinger et al., 2004). This is in addition to the time spent by learning scientists designing the pedagogy and dialog strategies. However, this is usually done only once for the topics in a subject area. As a result, this limits the AutoTutor to a small set of subject areas, restricting their impact. Nevertheless, there are many works in the space of rule-based ITS (Paladines and Ramirez, 2020), which have been reasonably successful in their domains.
@@ 3. Design Principles
The optimal tutoring support differs between the two examples.
@@ 3.3. Scaffolding the LLM
While modern LLMs fine-tuned to follow natural language instructions are quite good at doing so, they can at times struggle with more complex tasks. Because of the probabilistic nature of LLMs, reattempting the same problem can lead to different answers. To take the best advantage of these facts, we want to have automated checks for every LLM turn in the strategy space, so calls can be repeated until some checks have been met. For example, D3 could require that the numerical answer ( for A1) is included in the utterance, while P2 could require the missing keywords to be absent. These checks can take different forms depending on the domain, like limiting the length of an utterance or avoiding the use of more complex language.
3.4. The Trade-off between Reliability and Contextual Awareness
Once we have prompts for each utterance and relevant checks, we can query an LLM to generate the tutor utterances. Here we have a choice on when the generation is done. One option is to generate all possible utterances at design time, and storing them into the state space. This would be similar to what traditional ITSs would do. Doing this not only allows the tutor to run without an active internet connection, but having an exhaustive set of utterances stored allows stakeholders to trust the tutor without necessarily trusting the LLM used to generate the utterances. This is possible because one can reliably say that there would be no “new” errors or hallucinations since the LLM is not used after the design stage.
Despite these advantages, pre-generating means that all possible scenarios that can arise during the conversation need to be anticipated in advance, which is extremely hard to achieve. Therefore, such a model is likely to feel somewhat oblivious of the context of the conversation, which might reduce student engagement, which in turn, may reduce learning gains (Martin and Bolliger, 2018; Bryson and Hand, 2007; Parsons and Taylor, 2011) The alternative to this, is to generate the utterances during the conversation. This allows us to include recent utterances from both the tutor and the student into the prompt, thereby generating a more context-aware response. Due to changing context, it is also possible to allow multiple utterances for a single state, allowing for more fine-grained feedback. Doing this, however, leaves the LLM open to hallucinations and jailbreaks as discussed earlier.
The choice we make depends of factors like how good we believe the backend LLMs are, how much the policymakers trust them, how complex is the problem etc. We do not believe that any of these options is better than the other, and the best option might be a combination of both.
@@ 4. MWPTutor
We now detail the design process of our system MWPTutor, an instantiation of the above principles. MWPTutor is adapted to help students solve math word problems. The scenario, originally defined by MathDial (Macina et al., 2023), involves a student approaching a tutor after unsuccessfully attempting to solve a math word problem. The student’s incorrect solution is provided as a part of the input. GSM8k (Cobbe et al., 2021), the source for MathDial’s problems includes a single step-by-step solution to each problem, which can be seen as a fulfillment of the solution step space. However, to not become constrained to the availability of such solutions, and to allow diverse solution paths, we restrict MWPTutor to only the final numerical answer to the problem, and generate the solutions ourselves. Further, to highlight the tradeoff mentioned in section 3.4, we shall design two slightly different versions MWPTutor, which is allowed to make live calls to the LLM, and MWPTutor which is not. Note that unless otherwise specified, designs mentioned in this section apply to both versions.
@@ 4.3. Pedagogical Strategy
| Pump | asks the student to propose the next step. If the student proposes the step correctly, we ask them to solve it. |
| Hint | is only triggered if the student fails to propose a valid next step, or fails to execute the step they propose. It essentially gives out the step and asks the student for the calculation. This utterance should take the form of a question, and the only way to correctly answer the question should be to successfully execute the step. |
| Prompt | is triggered if the student fails to execute the step after the hint. This should push the student closer to the answer, without actually revealing the answer. |
| Assertion | is triggered if the student fails to answer even after the prompt. This should simply give out the answer, and have the student acknowledge it. |
In our running example, hints, prompts, and assertions correspond to D1, D2, and D3 respectively. The pump is not included.
@@
For MWPTutor, we allow up to prompts before moving to assertion. For both hints and prompts, the LLM is given the steps the student has correctly solved so far and the last utterances in the dialogue ( each for the student and the tutor) in addition to the solution step and strategy (see Section A.3 for details)
@@ 4.4. Answer and Path Detection
To execute the pedagogical strategies described above, we need to detect if the student has given a correct answer. Further, the hint state should only trigger if the student fails to propose a correct follow-up. For this, we need to have a way to match paths. We run into the same matching problems described in Section 4.2, and in case of MWPTutor, resort to similar basic techniques. The correct answer detection algorithm is shown in Algorithm 2. For states other than pump, it takes a single reference step and checks if its RHS is present in the student utterance. For the pump state, all possible followups of the last step are passed as reference steps. If it can find the RHS of any of these in the student utterance, it marks that as the next step and proceeds to the next step. If not, it considers a path match to have occurred if every operand on the left-hand side of some step is present. If multiple steps fulfill this, or no step does, we pick the one with the best BertScore alignment to the student utterance.
Since MWPTutor is allowed to make live calls to the LLM, we make use of it for both correct answer detection, and picking the best path to take when in the “pump” stage. For correct answer detection, we use the RHS match as an additional criterion, while for picking the best path, we rely entirely on the LLM’s judgement. The fact that MWPTutor currently ignores all non-numeric content in the student response is one of its major limitations, as it can miss a large number of “correct” utterances from the student, but we defer the resolution of this to future work.
@@ 4.5. Putting it All Together
Figure 3 summarises the overall state space of MWPTutor. It can be seen as a simplified version of the state space of AutoTutor (see Figure 1 in (Cai et al., 2019)). The expectation matching is implicitly carried out by the solution aligner and the pump state. To obtain to full finite state transducer structure, one must roll out the states along the solution space. Excluding the unlikely case where the student fails to repeat the answer after the assertion state, the overall conversation should not last more than utterance pairs for MWPTutor, and for MWPTutor, where is the number of steps.
@@ 5.1. Setup
At the time of testing, GPT4555gpt-4-1106-preview queried in January 2024 was considered to be one of the best-performing LLMs, so we use it both as a backend for the two MWPTutor’s and as the baseline single prompt LLM. To best align the single prompt GPT4 with the MWPTutor doctrine, we ask it to roleplay a tutor, setting the primary objective of getting to the correct answer as the goal of the conversation, and the secondary objective of not revealing the answer as a personality trait. Further, to test the potential of an open-source LLM as a backend, we have a second version of MWPTutor with a MetaMath-13B (Yu et al., 2023). MetaMath is a version of Meta AI’s open source LLM LLaMa666llama.meta.com fine-tuned to solve math word problems, and is, to the best of our knowledge, the best open LLM for solving math word problems. However, since MWPTutor would require the Backend to run throughout the conversation, which is unsustainable for a locally-run model, we only use this with MWPTutor. We shall refer to the live model as MWPTutor, the cached models as MWPTutor and MWPTutor based on the backend, and the single prompt model simply as GPT4 hereafter.
@@ 5.2. Conversation Evaluation
We evaluate both success and telling for the MultiArith dataset by directly looking for the answer in student and tutor utterances respectively, and calculate the overall scores based on that. The top half of Table 1 summarises our results. All models are able to transfer their near-perfect solving scores to near-perfect tutoring scores. Both GPT4 and MWPTutor have no telling, while the telling rate of the MWPTutor models is also almost negligble. All versions of MWPTutor have a perfect success rate, with GPT4 not far behind. These results demonstrate that all our models are aligned with our doctrine, and are able to follow it well on a simpler dataset. It must be noted that the metric is noisy since it only requires the number to match, with no restriction on the context, e.g. in the problem “How many groups of can be made out of boys and girls?” if the tutor mentions there are boys, we would count it as telling as the final answer happens to be . This means that the real rates could be somewhat lower for all models, and the difference between them is not significant. Nevertheless, such errors are unlikely to be frequent, and all models work reasonably well on these problems.
| MWPTutor | MWPT. | GPT4 | |||
| LLaMa | GPT4 | GPT4 | |||
| MultiArith | Success Rate | 1.00 | 1.00 | 1.00 | 0.98 |
| Telling Rate | 0.01 | 0.01 | 0.00 | 0.00 | |
| Adjusted Success | 0.99 | 0.99 | 1.00 | 0.98 | |
| Tutoring Score | 0.99 | 0.99 | 1.00 | 0.98 | |
| MathDial | Success Rate | 0.91 | 0.96 | 1.00 | 0.82 |
| Telling Rate | 0.33 | 0.11 | 0.00 | 0.00 | |
| Adjusted Success | 0.58 | 0.84 | 1.00 | 0.82 | |
| Tutoring Score | 0.71 | 0.90 | 1.00 | 0.82 | |
@@
The bottom half of Table 1 summarises our result. Both GPT4 and MWPTutor maintain their 0% telling rate, but MWPTutor is the only model able to maintain its 100% success rate. This clearly establishes MWPTutor as the model of choice should live access to GPT4 not be an issue, as discussed in Section 3.4.
Of the cached models, MWPTutor gets more problems correct than GPT4, and more when adjusted for telling. This gives MWPTutor an advantage on GPT4 in terms of the final score. MWPTutor also gets a higher success rate than GPT4, which is somewhat surprising since its backend model is significantly weaker than GPT4 in the MWP Solver setting (Yu et al., 2023). This increased success rate however comes at a cost of a very high telling rate, giving it a significantly poorer tutoring score overall. Based on this we do not recommend the use of MetaMath as the backend model unless GPT4-like models need to be avoided for some external reason.
One thing worth mentioning here is that once problematic questions have been identified for either version of MWPTutor, one can have human annotators redo only those questions, thereby increasing the success rate, something that is not possible for GPT4. MWPTutor is somewhere in the middle, as we can only “fix” the solution step space.
@@ 5.3. Utterance Evaluation
A major shortcoming of AI-on-AI-tutoring is that the AI student may not be affected by the sometimes subpar tutor responses. To better understand the quality of utterances both in terms of them executing the pedagogical strategy and in term of their context awareness, we hired human evaluators who worked in the education sector and asked them to evaluate utterances made by MWPTutor and MWPTutor for MathDial. Because the pumps are generic and very similar to each other and assertions essentially boil down to reading out the answers, we focus only on the hints and prompts in our evaluation.
@@
Across the conversations, MWPTutor and the student model went over solution steps Of these went to the hint state, while reached the prompt state. In the case of MWPTutor, we had steps total, of which went to a prompt hint state, went to the prompt stage. of these would then use a second prompt and would use a third, giving us a total of prompts. We presented each of these utterances to human evaluators, with the past conversation, the original math word problem, and the current solution step as context. We boil down the requirements set on hints and prompts set in Section 4.3 to the following binary criteria which we asked the evaluators to evaluate the utterances on:
Grammatical: Is linguistically and grammatically correct.
Not Reveal: Does not reveal the final answer for the step.
@@
Relevant (hints only): Addresses the step and the only way to correctly answer the question is to complete the current step.
@@
Not Redundant (hints only): Is not redundant to what the student already said in the previous step ie it makes sense to ask this question after the previous response.
@@
Specific (prompts only): Is specific to the step, and not just a vague lead-on statement.
@@
Helpful (prompts only): Can help the student make progress in the current step, given the previous student’s utterance.
@@
The annotators were paid the Prolific suggested wage of 10 GBP/hour. We show the annotation user interface in Figure 6 and some example annotations in Tables 2, LABEL:, 3, LABEL:, 4, LABEL: and 5. Figures 4 and 5 summarise the responses from the evaluators. Almost all utterances are considered Grammatical by the evaluators, which is in line with the known grammatical prowess of GPT4. It is unclear why there is a slightly higher count of non-grammatical utterances in MWPTutor, but it could be potentially due to annotators finding the more eloquent responses of GPT4 harder to judge grammatically. Most of the utterances are also marked to be non-revealing, which is expected given that any proposed utterances that directly revealed the answer were assigned to be regenerated. MWPTutor performs slightly better here. The evaluators further seem to agree that most hints were relevant to the step and that most prompts are specific, implying that they do what is appropriate given the current solution step. Again, MWPTutor has an edge in this aspect. Of the two of its prompts found to be non-specific, one is a second prompt and one is a first prompt, so all third prompts are considered specific
The biggest potential weakness of MWPTutor comes in the form of its utterances making sense in the context of previous student utterances due to issues discussed in Section 4.4, which is where the final two metrics come in. In this regard, all prompts from MWPTutor are considered to be helpful by a majority of evaluators which is a great sign. Surprisingly, MWPTutor ends up giving 5 prompts that are not considered helpful. One reason why this happens could be the fact that there are more prompts allowed for each step, making individual prompts less helpful. Regardless, it is still a small number.
Hints, however, are found to be redundant in almost a third of cases for MWPTutor, showing that handling of student’s responses to pumps could be better. MWPTutor brings this number much lower, but still with redundant Hints, it is the second largest red bar in the graph, only behind its cached counterpart. It is clear that MWPTutor’s heuristic of matching the left-hand side is too demanding, and highly likely to ignore the student even if they propose a valid next step to the problem, which can break student immersion.
@@
It should be noted, however that of the steps did not proceed to the hint state for MWPTutor, implying more than half of the correct proposals by the student are being identified as such. Curiously, the number of non-redundant hints is almost the same for the two models. Further, in a vast majority of cases where a correct proposal is misidentified, there is disagreement between human evaluators indicating there might be some merit in the rejection, though this is also the case for a many of correct rejections.
@@
Overall these evaluations show that utterances generated for MWPTutor by the GPT4 backend are consistent with the strategy definitions, and the transition function uses them reasonably well. MWPTutor further increases the efficacy of the transition function. Both models, however, have scope for improvement.